OpenStudy (anonymous):

The circle graph shows what percentage of students at Rockford college are enrolled in various majors. Fine the ratio of science majors to business majors science major is: 9% business major is: 31%

jimthompson5910 (jim_thompson5910):

Say there are 100 students total. How many are science majors? How many are business majors?

OpenStudy (anonymous):

business major is 31% science is 9%

OpenStudy (anonymous):

\[\frac{ 9 }{ 31}=\frac{ x }{ 100 }\]

OpenStudy (anonymous):

no ?

jimthompson5910 (jim_thompson5910):

not quite

jimthompson5910 (jim_thompson5910):

what is 9% of 100?

OpenStudy (anonymous):

9 ?

jimthompson5910 (jim_thompson5910):

and what is 31% of 100

OpenStudy (anonymous):

31

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

So the ratio of the two is 9:31

OpenStudy (anonymous):

wow thanks

OpenStudy (anonymous):

y=\frac{ 1 }{ 5}x-4 Complete the ordered pairs (5, ) (0, )

OpenStudy (anonymous):

what about this

jimthompson5910 (jim_thompson5910):

all ordered pairs look like this (x,y) so when they say for instance (2,3) it means (x,y) = (2,3) x = 2 y = 3

jimthompson5910 (jim_thompson5910):

so (5, ) means x = 5 y = unknown

jimthompson5910 (jim_thompson5910):

you then plug x = 5 into \[\Large y=\frac{ 1 }{ 5}x-4\] and evaluate to find y

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

is it 21 ?

jimthompson5910 (jim_thompson5910):

no

OpenStudy (anonymous):

1/5(5)-4

OpenStudy (anonymous):

im confused

jimthompson5910 (jim_thompson5910):

(1/5)*5 - 4 = ??

OpenStudy (anonymous):

3

jimthompson5910 (jim_thompson5910):

close

OpenStudy (anonymous):

-3

jimthompson5910 (jim_thompson5910):

better

jimthompson5910 (jim_thompson5910):

when x = 5, y = -3

OpenStudy (anonymous):

thats correct ?

jimthompson5910 (jim_thompson5910):

meaning we have this ordered pair: (5, -3)

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

theres 2 im suppose to have

OpenStudy (anonymous):

5,-3 and ( 0, )

jimthompson5910 (jim_thompson5910):

plug in x = 0 to get y = ??

OpenStudy (anonymous):

will it be 4 ?

jimthompson5910 (jim_thompson5910):

nope

OpenStudy (anonymous):

-4

jimthompson5910 (jim_thompson5910):

so another ordered pair is (0,-4)

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

np